As shown above, a classic deck of cards is made up of 52 cards, 26 are black, 26 are red. Each color is split into two suits of 13 cards each (clubs and spades are black and hearts and diamonds are red). Each suit is split into 13 individual cards (Ace, 2-10, Jack, Queen, and King).
If you select a card at random, what is the probability of getting: (Round to 4 decimal places where possible)
a) A 3 of Diamonds?

b) A Diamond?

c) A number smaller than 6 (counting the ace as a 1)?

Answer by mananth(16946)


a) A 3 of Diamonds?
only 1 favourable outcome out of 52 possible outcomes = 1/52
b) A Diamond?
There are 13 diamond cards in a suit 13 favourable outcomes 13/52 =1/4
c) A number smaller than 6 (counting the ace as a 1)?
5 in each suit 4 suits
favoirable outcomes = 5*4 = 20
= 20/52 = 10/26
a) Probability of a 3 of Diamonds: 1/52 ≈ 0.0192
b) Probability of a Diamond: 1/4 = 0.25
c) Probability of a number smaller than 6: 20/52 ≈ 0.3846